At sea level, the boiling point of water is 100 degrees Celsius. At an altitude of 2 km, the boiling point is 93 degrees Celsius.
A. Write a linear function for the boiling point of water of the altitude above sea level.
B. Use your function to predict the boiling point of water on top of Mount Everest, which is approximately 8.85 km above sea level. Round to nearest degree.

Answer by fractalier(6550)

Call the boiling point temperature, T. It is a function of altitude, a.
We can then write a linear function as
T(a) = 100 - 3.5a
by using the two points (0, 100) and (2, 93).
For a = 8.85,
T(8.85) = 100 - (3.5)(8.85) = 69.025 degrees C